OpenStudy (anonymous):

Determine the rate of change for the equation 1/6 y + 3/8 x = 24

OpenStudy (anonymous):

[-1/(6y^2)]*dy/dx-[3/(8x^2)]=0 dy/dx={3/(8x^2)]*(-6y^2) dy/dx=-9y^2/4x^2

OpenStudy (anonymous):

What does the d represent?

OpenStudy (anonymous):

dy represents infinitesimal change in y and dy/dx represents change in y with respect to x instantaneously. dy/dx is the rate change of y w.r.t x

OpenStudy (anonymous):

so do I follow the equation for finding the slope y=mx+b ?

OpenStudy (anonymous):

Yes. Get it into the form y = mx + b once it is in that form, the slope of the line is m

OpenStudy (anonymous):

Ok can you help me with that. The fractions are loosing me here.

OpenStudy (anonymous):

@keegq 1/6 y + 3/8 x = 24 \[\frac{ 1 }{ 6 } \frac{ dy }{ dx } + \frac{ 3 }{ 8 } = 0\]\[\frac{ dy }{ dx } = \frac{ -3 }{ 8 } *6\]\[\frac{ dy }{ dx } = \frac{ -9 }{ 8 }\] I don't know what you did up there but it ain't right.

OpenStudy (anonymous):

Thank you for your help!!

OpenStudy (anonymous):

oops, i didnt mean to reply to you keegq. I meant to do @mayankD

OpenStudy (anonymous):

I knew what you meant

OpenStudy (anonymous):

So the rate of change is -9/8

OpenStudy (anonymous):

Dude are y and x in the numerator or the denominator?